### v2

@@ 4 Proximal point method for convex optimization
In this section, we will utilize the regularity conditions in Section 3 to derive fast linear convergence guarantees of the classical proximal point method (PPM) [43] for convex (potentially nonsmooth) optimization and generalize it to weakly convex settings. PPM is a conceptually simple algorithm, which has been historically used for guiding algorithm design and analysis, such as proximal bundle methods [44] and augmented Lagrangian methods [43]. It has recently found increasing applications in modern machine learning; see [32].
@@ 4.1 Proximal point method
$f^{\star}=\min_{x}\;f(x),\vspace{-1mm}$ (10)
where $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is a proper closed convex function. Note that 10 is also an abstract model for constrained optimization, since given a closed convex set $X$, we can define $\bar{f}(x)=f(x)$ if $x\in X$, and $\bar{f}(x)=+\infty$ otherwise. Let $S=\operatorname*{\arg\!\min}_{x}\,f(x)$. We define the proximal mapping as
@@
where $\{c_{k}\}_{k\geq 0}$ is a sequence of positive real numbers. The quadratic term in 11 makes the objective function strongly convex and always admits a unique solution. The iterates 12 are thus well-defined.
The convergence of PPM 12 for (nonsmooth) convex optimization has been studied since the 1970s [24]. The sublinear convergence is relatively easy to establish, and many different assumptions exist for linear convergences of 12; see [24, 33, 16, 25, 8]. However, as we will highlight later, some assumptions are restrictive and the corresponding proofs are sophisticated and nontransparent. We aim to provide simple proofs under the general regularity conditions in Section 3.
@@
Under a very general setup, the PPM 12 converges at a sublinear rate for cost value gaps, and the iterates converge asymptotically, as summarized in Theorem 4.1. This result is classical [45, Theorem 2.1], and a new bound with an improved constant 4 is also available in [46, Theorem 4.1] using the performance estimation technique.
Theorem 4.1 (Sublinear convergence [45, Theorem 2.1]).
Let $f\!:\!\mathbb{R}^{n}\!\to\!\overline{\mathbb{R}}$ be a proper closed convex function with $S=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x)$ and $S\neq\emptyset$. Then, the iterates 12 with a positive sequence $\{c_{k}\}_{k\geq 0}$ satisfy
@@
The proof of 13 is immediate from a telescope sum via the following one-step improvement:
@@
This fact is not difficult to establish. Note that choosing any constant step size $c_{k}=c>0$ in 13 directly implies the common sublinear rate $\mathcal{O}(1/k)$. In Theorem 4.1, $f$ does not need to be $L$-smooth, and it can also be non-differentiable. Thus, the guarantees in Theorem 4.1 are much stronger than those by (sub)gradient methods. This is because the proximal mapping 11 is a stronger oracle than simple (sub)gradient updates.
@@
Let $f\!:\!\mathbb{R}^{n}\!\to\!\overline{\mathbb{R}}$ be a proper closed convex function with $S=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x)$, $S\neq\emptyset$, and $\nu>0$. Suppose $f$ satisfies PL (or EB, RSI, QG) over the sublevel set $[f\leq f^{\star}+\nu]$. Then, for all $k\geq k_{0}=\frac{\mathrm{dist}^{2}(x_{0},S)}{2\nu\inf_{k\geq 0}c_{k}}$ steps, the iterates 12 with a positive sequence $\{c_{k}\}_{k\geq 0}$ bounded away from zero enjoy linear convergence rates, i.e.,
@@
The sublinear convergence in Theorem 4.1 ensures that the iterate $x_{k}$ reaches $[f\leq f^{\star}+\nu]$ after at most $k_{0}$ iterations. Once $x_{k}$ is within $[f\leq f^{\star}+\nu]$, all the properties EB, PL, RSI, and QG are equivalent by Theorem 3.1, and the next point $x_{k+1}$ also satisfies $x_{k+1}\in[f\leq f^{\star}+\nu]$ as $f(x_{k+1})\leq f(x_{k+1})+\frac{1}{2c_{k}}\|x_{k+1}-x_{k}\|^{2}\leq f(x_{k})$ by the update rule 12. For the analysis below, we assume $x_{k}\in[f\leq f^{\star}+\nu]$.
We next show that PL gives a simple proof of 15a, and QG together with EB leads to a clean proof of 15b. Recall that the optimality condition of 12 directly implies
$-(x_{k+1}-x_{k})/c_{k}\in\partial f(x_{k+1}).$ (16)
@@
where $(a)$ applies the fact that $x_{k}$ is a suboptimal solution to 12, $(b)$ comes from the optimality 16, and $(c)$ applies PL. Re-arranging and subtracting $f^{*}$ from both sides of 17 lead to the desired linear convergence in 15a.
We next use QG to prove 15b with coefficient $\theta_{k}\leq 1/\sqrt{c_{k}\mu_{\mathrm{q}}+1}$. By definition, we have $f(\Pi_{S}(x_{k}))=f^{\star}$ and $\|\Pi_{S}(x_{k})-x_{k}\|^{2}=\mathrm{dist}^{2}(x_{k},S)$. Since $f+\frac{1}{2c_{k}}\|\cdot-x_{k}\|^{2}$ is $1/c_{k}$ strongly convex, its first-order lower bound at $x_{k+1}$ is
@@
where we also applied the fact that $x_{k+1}$ minimizes 12 so 0 is a subgradient. From 18, we drop the positive term $\|x_{k+1}-x_{k}\|^{2}$ and use the fact that $\|\Pi_{S}(x_{k})-x_{k+1}\|\geq\mathrm{dist}(x_{k+1},S)$, leading to
@@
Simple arguments based on EB can establish 15b with coefficient $\theta_{k}\leq 1/(\mu_{\mathrm{e}}^{2}/c_{k}^{2}+1)^{1/2}$. We provide some details in Section C.1.∎
Two nice features of Theorem 4.2 are 1) the simplicity of its proofs and 2) the generality of its conditions. Indeed, the proof of 15a is simple via PL, and the proof of 15b is clean via QG and EB, which are simpler than typical proofs. In addition, the regularity conditions are weaker than [24] and [33]. Linear convergence for $\mathrm{dist}(x_{k},S)$ was first established in [24] with one restrictive assumption: the inverse of the subdifferential $(\partial f)^{-1}$ is locally Lipschitz at 00, which requires a unique optimal solution, i.e., $S$ is a singleton. The uniqueness assumption is lifted in [33], which even allows an unbounded solution set. More recently, this assumption is further relaxed to $\partial f$ being metrically subregular in [25] and [16]. It is known that for convex functions, $\partial f$ is metrically subregular if and only if $f$ satisfies quadratic growth (cf. Theorem 3.1). Indeed, our proof in Theorem 4.2 is based on quadratic growth, which is a more intuitive geometrical property. Our main idea in the proof above is partially inspired by a recent result for the linear convergence of the spectral bundle method in [15, 14].
We note that in the proof of Theorem 4.2, once the iterate is within the PL/QG region, the proof idea follows simply on the existence of the optimality condition 16. Building upon this observation, we extend Theorem 4.2 to the class of weakly convex functions when a proper initialization is given.
@@
Let $f\!:\!\mathbb{R}^{n}\!\to\!\overline{\mathbb{R}}$ be a proper closed $\rho$-weakly convex function with $S=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x)$ and $S\neq\emptyset$. Suppose $f$ satisfies QG with $\mu_{\mathrm{q}}>\rho$ over the sublevel set $[f\leq f^{\star}+\nu]$ with $\nu>0$. Let $\beta=\mu_{\mathrm{q}}-\rho>0$. The iterates generated by the PPM 12 with any positive sequence $\{c_{k}\}_{k\geq 0}$ satisfying $1/c_{k}>\rho$ for all $k\geq 0$ and an initialization $x_{0}\in[f\leq f^{\star}+\nu]$ satisfy
@@
$\omega_{k}=\frac{1}{1+\mu_{\mathrm{p}}c_{k}}<1\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \theta_{k}=\min\mathopen{}\mathclose{{}\left\{\frac{1}{\sqrt{c_{k}\beta+1}},\frac{1}{\sqrt{c_{k}^{2}/\mu_{\mathrm{e}}^{2}+1}}}\right\}<1,$
@@
We first note that choosing $\{c_{k}\}_{k\geq 0}$ such that $1/c_{k}>\rho,\forall k\geq 0$, guarantees the existence of the optimality condition $-(x_{k+1}-x_{k})/c_{k}\in\hat{\partial}f(x_{k+1})$ similar to 16. With the existence of the optimality condition, the proof follows similar arguments in the proof of Theorem 4.2. Specifically, the proof of the linear decrease in the cost value gap is identical to that in Theorem 4.2. As $\mu_{\mathrm{q}}>\rho$, EB, RSI, and PL hold by Theorem 3.1. For the proof of the linear decrease in the distance, the coefficient $(c_{k}^{2}/\mu_{\mathrm{e}}^{2}+1)^{-1/2}$ follows the same argument as in the convex case since $c_{k}$ is chosen such that $1/c_{k}>\rho$ so the function $f+\frac{1}{2c_{k}}\|\cdot-x_{k}\|^{2}$ is a convex function plus a quadratic term and the proximal mapping satisfied the key inequality in Lemma C.1 in Section C.1; For the coefficient $(c_{k}\beta+1)^{-1/2}$, the argument is essentially the same as the convex case with a small modification that the function $f+\frac{1}{2c_{k}}\|\cdot-x_{k}\|^{2}$ is $(\frac{1}{c_{k}}-\rho)$-strongly convex. ∎
@@
Example 4.1 (PPM for a weakly convex function).
Consider the 2-weakly convex function $f$ in Example 3.1. It satisfies QG and $26=\mu_{\mathrm{q}}>\rho=$ 2 globally. Therefore, any initial point $x_{0}\in\mathbb{R}$ with a positive sequence $\{c_{k}\}_{k\geq 0}$ such that $1/c_{k}>\rho$ for all $k\geq 0$ guarantees the linear convergence. Let $c_{k}=1/6$ for all $k\geq 0$. For any $x_{k}\in\mathbb{R}$, from a simple calculation, the PPM update 12 follows the formula
@@
$v_{1}=\begin{cases}-0.5,&\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \frac{3}{2}x_{k}\geq-0.5\\ \frac{3}{2}x_{k},&\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ -1<\frac{3}{2}x_{k}<-0.5\\ -1,&\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \frac{3}{2}x_{k}\leq-1\end{cases}$
@@
$v_{2}=\begin{cases}\operatorname*{\arg\!\min}_{x\in\{-1,-0.5\}}f(x)+3(x-x_{k})^{2},&\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ -1<\frac{x_{k}-1}{2}<-0.5\\ \frac{x_{k}-1}{2},&\leavevmode\nobreak\ \text{otherwise}.\end{cases}$
We present the numerical result with the initial point $x_{0}=-2$ in Figure 2. It clearly shows the linear convergence in both the distance to the solution set and the cost value gap.
Figure 2: Linear convergences of the distance to the solution set (left) and cost value gaps (right) for a $2$-weakly convex function in Example 4.1.
@@ 5 Inexact proximal point method (iPPM) and its convergence
In Section 4, each subproblem 11 is solved exactly. This may not be practical since one still needs an iterative solver to solve 11, where stopping criteria naturally introduce errors. We here discuss an inexact version of PPM (iPPM, [24]) where the subproblem 12 is solved inexactly. The regularity conditions in Theorem 3.1 also allow us to establish linear convergences of iPPM.
@@
We replace the exact update 12 with an inexact update
@@
Two classical criteria suggested in Rockafellar’s seminal work [24] are
@@
The inexact update 20 with A or B is called iPPM. The two criteria are not directly implementable as the value of $\text{prox}_{c_{k}f}(x_{k})$ is unknown. As discussed in [24, Proposition 3], two implementable alternatives that imply A and B, respectively, are
@@
The seminal work by Rockafellar [24] has established the asymptotic convergence of iterates for iPPM under a general setup. We state the results below whose proof is technically involved.
Theorem 5.1 (Asymptotic convergence of iterates [24, Theorem 1]).
Consider a proper closed convex function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$. Let $\{x_{k}\}_{k\geq 0}$ be any sequence by 20 under A with a positive sequence $\{c_{k}\}_{k\geq 0}$ bounded away from zero. Then, we have 1) the sequence $\{x_{k}\}_{k\geq 0}$ is bounded if and only if there exists a solution to $0\in\partial f(x)$, i.e., $S\neq\emptyset$; 2) if $S\neq\emptyset$, the whole sequence $\{x_{k}\}_{k\geq 0}$ converges to an optimal point $x_{\infty}\in S$, i.e., $\lim_{k\to\infty}x_{k}=x_{\infty}$.
We next establish a sublinear convergence of iPPM for cost value gaps. Our simple proof is based on the boundedness of the iterates from Theorem 5.1 and a recent idea in [47, Theorem 3]. We provide its proof details in Appendix D.
@@
Let $f:\mathbb{R}^{n}\!\to\!\overline{\mathbb{R}}$ be a proper closed convex function with $S=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x)$ and $S\neq\emptyset$. The iterates 20 under $\text{A}^{\prime}$ with a positive sequence $\{c_{k}\}_{k\geq 0}$ bounded away from zero converge to $x^{\star}\in S$ asymptotically, and the cost value gaps satisfy
@@
Note that Theorems 5.2 and 5.1 can be viewed as the convergence counterpart for iPPM of Theorem 4.1 with two major differences: 1) Theorem 5.2 deals with the best iterate, unlike the last iterate in Theorem 4.1 (the guarantee for the average $\bar{x}_{k}=\frac{1}{k}\sum_{j=1}^{k}x_{j}$ or weighted average $\tilde{x}_{k}=({\sum_{j=0}^{k-1}c_{j}x_{j+1}})/({\sum_{j=0}^{k-1}c_{j}})$ is also straightforward; see remark D.1 in the appendix); 2) the convergence of cost values in Theorem 5.2 relies on the boundedness of iterates in Theorem 5.1 whose proof is technically involved [24, Theorem 1], while the convergence of iterates of exact PPM is established from the sublinear convergence of cost values in Theorem 4.1.
Similar to Theorem 4.2, linear convergence of iPPM can also be deduced when popular regularity conditions are assumed. However, the definitions of SC, RSI, EB, PL, and QG in Section 3 are stated over a sublevel set $[f\leq f^{\star}+\nu]$ with $\nu>0$, which is convenient to estimate the iteration number to enter a certain region, such as the constant $k_{0}$ in Theorem 4.2; but those definitions become too stringent in the analysis. Indeed, note that the proof in Theorem 4.2 requires the next point $x_{k+1}$ remains inside the QG or PL region. As the stopping criteria A and B allow infeasible points, it is possible that $x_{k+1}$ is infeasible (i.e., $f(x_{k+1})=+\infty$), making the same proof fail. However, the iterates are still approaching an optimal solution and the inequality $\mu_{\mathrm{q}}/2\cdot\mathrm{dist}(x_{k+1},S)\leq f(x_{k+1})-f^{\star}$ with any $\mu_{\mathrm{q}}>0$ still holds automatically.
@@
Notice that we can also redefine RSI, EB, and PL using a neighborhood (similar to 21) and show the equivalence between them. However, we note that the neighborhood $\mathcal{U}$ for different regularity conditions may be different. For clarity, we present our final technical result that shows the linear convergence of iPPM under 21, which is the counterpart of Theorem 4.2.
@@
Let $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be a proper closed convex function with $S=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x)$ and $S\neq\emptyset$. Suppose $f$ satisfies 21 at all $x^{\star}\in S$. Let $\{x_{k}\}_{k\geq 0}$ be any sequence generated by iPPM 20 under A and B with a positive sequence $\{c_{k}\}_{k\geq 0}$ bounded away from zero. Then, there exist a nonnegative $\theta_{k}<1$ and a large $\bar{k}>0$ such that

### v3

@@ 4 Proximal point method for convex optimization
In this section, we will utilize the regularity conditions in Section 3 to derive fast linear convergence guarantees of the classical proximal point method (PPM) [rockafellar1976augmented] for convex (potentially nonsmooth) optimization and generalize it to weakly convex settings. PPM is a conceptually simple algorithm, which has been historically used for guiding algorithm design and analysis, such as proximal bundle methods [lemarechal1981bundle] and augmented Lagrangian methods [rockafellar1976augmented]. It has recently found increasing applications in modern machine learning; see [drusvyatskiy2017proximal].
@@ 4.1 Proximal point method
$f^{\star}=\min_{x}\;f(x),\vskip-2.84526pt$ (13)
where $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is a proper closed convex function. Note that 13 is also an abstract model for constrained optimization, since given a closed convex set $X$, we can define $\bar{f}(x)=f(x)$ if $x\in X$, and $\bar{f}(x)=+\infty$ otherwise. Let $S=\operatorname*{\arg\!\min}_{x}\,f(x)$. We define the proximal mapping as
@@
where $\{c_{k}\}_{k\geq 0}$ is a sequence of positive real numbers. The quadratic term in 14 makes the objective function strongly convex and always admits a unique solution. The iterates 15 are thus well-defined.
The convergence of PPM 15 for (nonsmooth) convex optimization has been studied since the 1970s [rockafellar1976monotone]. The sublinear convergence is relatively easy to establish, and many different assumptions exist for linear convergences of 15; see [rockafellar1976monotone, luque1984asymptotic, leventhal2009metric, cui2016asymptotic, drusvyatskiy2018error]. However, as we will highlight later, some assumptions are restrictive and the corresponding proofs are sophisticated and nontransparent. We aim to provide simple proofs under the general regularity conditions in Section 3.
@@
Under a very general setup, the PPM 15 converges at a sublinear rate for cost value gaps, and the iterates converge asymptotically, as summarized in Theorem 4.1. This result is classical [guler1991convergence, Theorem 2.1], and a new bound with an improved constant 4 is also available in [taylor2017exact, Theorem 4.1] using the performance estimation technique.
Theorem 4.1 (Sublinear convergence [guler1991convergence, Theorem 2.1]).
Let $f\!:\!\mathbb{R}^{n}\!\to\!\overline{\mathbb{R}}$ be a proper closed convex function with $S=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x)$ and $S\neq\emptyset$. Then, the iterates 15 with a positive sequence $\{c_{k}\}_{k\geq 0}$ satisfy
@@
The proof of 16 is immediate from a telescope sum via the following one-step improvement:
@@
This fact is not difficult to establish. Note that choosing any constant step size $c_{k}=c>0$ in 16 directly implies the common sublinear rate $\mathcal{O}(1/k)$. In Theorem 4.1, $f$ does not need to be $L$-smooth, and it can also be non-differentiable. Thus, the guarantees in Theorem 4.1 are much stronger than those by (sub)gradient methods. This is because the proximal mapping 14 is a stronger oracle than simple (sub)gradient updates.
@@
Let $f\!:\!\mathbb{R}^{n}\!\to\!\overline{\mathbb{R}}$ be a proper closed convex function with $\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x), SS\neq\emptyset$, and $\nu>0$. Suppose $f$ satisfies PL (or EB, RSI, QG) over the sublevel set $f\leq f^{\star}+\nu]$. Then, for all $k\geq k_{0}=\frac{\mathrm{dist}^{2}(x_{0},S)}{2\nu\inf_{k\geq 0}c_{k}}$ steps, the iterates 15 with a positive sequence $\{c_{k}\}_{k\geq 0}$ bounded away from zero enjoy linear convergence rates, i.e.,
@@
The sublinear convergence in Theorem 4.1 ensures that the iterate $x_{k}$ reaches $f\leq f^{\star}+\nu]$ after at most $k_{0}$ iterations. Once $x_{k}$ is within $f\leq f^{\star}+\nu]$, all the properties EB, PL, RSI, and QG are equivalent by Theorem 3.1, and the next point $x_{k+1}$ also satisfies $x_{k+1}\in[f\leq f^{\star}+\nu]$ as $f(x_{k+1})\leq f(x_{k+1})+\frac{1}{2c_{k}}\|x_{k+1}-x_{k}\|^{2}\leq f(x_{k})$ by the update rule 15. For the analysis below, we assume $x_{k}\in[f\leq f^{\star}+\nu]$.
We next show that PL gives a simple proof of 18a, and QG together with EB leads to a clean proof of 18b. Recall that the optimality condition of 15 directly implies
$(x_{k+1}-x_{k})/c_{k}\in\partial f(x_{k+1}).$ (19)
@@
where $(a)$ applies the fact that $x_{k}$ is a suboptimal solution to 15, $(b)$ comes from the optimality 19, and $(c)$ applies PL. Re-arranging and subtracting $f^{*}$ from both sides of 20 lead to the desired linear convergence in 18a.
We next use QG to prove 18b with coefficient $\theta_{k}\leq 1/\sqrt{c_{k}\mu_{\mathrm{q}}+1}$. By definition, we have $f(\Pi_{S}(x_{k}))=f^{\star}$ and $\|\Pi_{S}(x_{k})-x_{k}\|^{2}=\mathrm{dist}^{2}(x_{k},S)$. Since $f+\frac{1}{2c_{k}}\|\cdot-x_{k}\|^{2}$ is $1/c_{k}$ strongly convex, its first-order lower bound at $x_{k+1}$ is
@@
where we also applied the fact that $x_{k+1}$ minimizes 15 so 0 is a subgradient. From 21, we drop the positive term $\|x_{k+1}-x_{k}\|^{2}$ and use the fact that $\|\Pi_{S}(x_{k})-x_{k+1}\|\geq\mathrm{dist}(x_{k+1},S)$, leading to
@@
Simple arguments based on EB can establish 18b with coefficient $\theta_{k}\leq 1/(\mu_{\mathrm{e}}^{2}/c_{k}^{2}+1)^{1/2}$. We provide some details in Section C.1. ∎
Two nice features of Theorem 4.2 are 1) the simplicity of its proofs and 2) the generality of its conditions. Indeed, the proof of 18a is simple via PL, and the proof of 18b is clean via QG and EB, which are simpler than typical proofs. In addition, the regularity conditions are weaker than [rockafellar1976monotone] and [luque1984asymptotic]. Linear convergence for $\mathrm{dist}(x_{k},S)$ was first established in [rockafellar1976monotone] with one restrictive assumption: the inverse of the subdifferential $(\partial f)^{-1}$ is locally Lipschitz at 0, which requires a unique optimal solution, i.e., $S$ is a singleton. The uniqueness assumption is lifted in [luque1984asymptotic], which even allows an unbounded solution set. More recently, this assumption is further relaxed to $\partial f$ being metrically subregular in [cui2016asymptotic] and [leventhal2009metric]. It is known that for convex functions, $\partial f$ is metrically subregular if and only if $f$ satisfies quadratic growth (cf. Theorem 3.1). Indeed, our proof in Theorem 4.2 is based on quadratic growth, which is a more intuitive geometrical property. Our main idea in the proof above is partially inspired by a recent result for the linear convergence of the spectral bundle method in [ding2023revisiting, liao2023overview].
We note that in the proof of Theorem 4.2, once the iterate is within the PL/QG region, the proof idea follows simply on the existence of the optimality condition 19. Building upon this observation, we extend Theorem 4.2 to the class of weakly convex functions when a proper initialization is given.
@@
Let $f\!:\!\mathbb{R}^{n}\!\to\!\overline{\mathbb{R}}$ be a proper closed $\rho$-weakly convex function with $S=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x)$ and $S\neq\emptyset$. Suppose $f$ satisfies QG with $\mu_{\mathrm{q}}>\rho$ over the sublevel set $f\leq f^{\star}+\nu]$ with $\nu>0$. Let $\beta=\mu_{\mathrm{q}}-\rho>0$. The iterates generated by the PPM 15 with any positive sequence $\{c_{k}\}_{k\geq 0}$ satisfying $1/c_{k}>\rho$ for all $k\geq 0$ and an initialization $x_{0}\in[f\leq f^{\star}+\nu]$ satisfy
@@
$\omega_{k}=\frac{1}{1+\mu_{\mathrm{p}}c_{k}}<1\penalty 10000\ \text{and}\penalty 10000\ \theta_{k}=\min\mathopen{}\mathclose{{\left\{\frac{1}{\sqrt{c_{k}\beta+1}},\frac{1}{\sqrt{c_{k}^{2}/\mu_{\mathrm{e}}^{2}+1}}}}\right\}<1,$
@@
We first note that choosing $\{c_{k}\}_{k\geq 0}$ such that $1/c_{k}>\rho,\forall k\geq 0$, guarantees the existence of the optimality condition $(x_{k+1}-x_{k})/c_{k}\in\hat{\partial}f(x_{k+1})$ similar to 19. With the existence of the optimality condition, the proof follows similar arguments in the proof of Theorem 4.2. Specifically, the proof of the linear decrease in the cost value gap is identical to that in Theorem 4.2. As $\mu_{\mathrm{q}}>\rho$, EB, RSI, and PL hold by Theorem 3.1. For the proof of the linear decrease in the distance, the coefficient $(c_{k}^{2}/\mu_{\mathrm{e}}^{2}+1)^{-1/2}$ follows the same argument as in the convex case since $c_{k}$ is chosen such that $1/c_{k}>\rho$ so the function $f+\frac{1}{2c_{k}}\|\cdot-x_{k}\|^{2}$ is a convex function plus a quadratic term and the proximal mapping satisfied the key inequality in Lemma C.1 in Section C.1; For the coefficient $(c_{k}\beta+1)^{-1/2}$, the argument is essentially the same as the convex case with a small modification that the function $f+\frac{1}{2c_{k}}\|\cdot-x_{k}\|^{2}$ is $(\frac{1}{c_{k}}-\rho)$-strongly convex. ∎
@@
Example 2 (PPM for a weakly convex function).
Consider the 2-weakly convex function $f$ in Example 1. It satisfies QG and $6=\mu_{\mathrm{q}}>\rho=2$ globally. Therefore, any initial point $x_{0}\in\mathbb{R}$ with a positive sequence $\{c_{k}\}_{k\geq 0}$ such that $1/c_{k}>\rho$ for all $k\geq 0$ guarantees the linear convergence. Let $c_{k}=1/6$ for all $k\geq 0$. For any $x_{k}\in\mathbb{R}$, from a simple calculation, the PPM update 15 follows the formula
@@
$v_{1}=\begin{cases}-0.5,&\penalty 10000\ \text{if}\penalty 10000\ \frac{3}{2}x_{k}\geq-0.5\\ \frac{3}{2}x_{k},&\penalty 10000\ \text{if}\penalty 10000\ -1<\frac{3}{2}x_{k}<-0.5\\ -1,&\penalty 10000\ \text{if}\penalty 10000\ \frac{3}{2}x_{k}\leq-1\end{cases}$
@@
$v_{2}=\begin{cases}\operatorname*{\arg\!\min}_{x\in\{-1,-0.5\}}f(x)+3(x-x_{k})^{2},&\penalty 10000\ \text{if}\penalty 10000\ -1<\frac{x_{k}-1}{2}<-0.5\\ \frac{x_{k}-1}{2},&\penalty 10000\ \text{otherwise}.\end{cases}$
We present the numerical result with the initial point $x_{0}=-2$ in Figure 2. It shows the linear convergence in both the distance to the solution set and the cost value gap.
Figure 2: Linear convergences of the distance to the solution set (left) and cost value gaps (right) for a $2$-weakly convex function in Example 2.
@@ 5 Inexact proximal point method (iPPM) and its convergence
In Section 4, each subproblem 14 is solved exactly. This may not be practical since one still needs an iterative solver to solve 14, where stopping criteria naturally introduce errors. We here discuss an inexact version of PPM (iPPM, [rockafellar1976monotone]) where the subproblem 15 is solved inexactly. The regularity conditions in Theorem 3.1 also allow us to establish linear convergences of iPPM.
@@
We replace the exact update 15 with an inexact update
@@
Two classical criteria suggested in Rockafellar’s seminal work [rockafellar1976monotone] are
@@
The inexact update 23 with A or B is called iPPM. The two criteria are not implementable as the value of $\text{prox}_{c_{k}f}(x_{k})$ is unknown. As discussed in [rockafellar1976monotone, Proposition 3], two implementable alternatives that imply A and B, respectively, are
@@
The seminal work by Rockafellar [rockafellar1976monotone] has established the asymptotic convergence of iterates for iPPM under a general setup. We state the results below, whose proof is technically involved.
Theorem 5.1 (Asymptotic convergence of iterates [rockafellar1976monotone, Theorem 1]).
Consider a proper closed convex function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$. Let $\{x_{k}\}_{k\geq 0}$ be any sequence by 23 under A with a positive sequence $\{c_{k}\}_{k\geq 0}$ bounded away from zero. Then, we have 1) the sequence $\{x_{k}\}_{k\geq 0}$ is bounded if and only if there exists a solution to $0\in\partial f(x)$, i.e., $S\neq\emptyset$; 2) if $S\neq\emptyset$, the whole sequence $\{x_{k}\}_{k\geq 0}$ converges to an optimal point $x_{\infty}\in S$, i.e., $\lim_{k\to\infty}x_{k}=x_{\infty}$.
We next establish a sublinear convergence of iPPM for cost value gaps. Our simple proof is based on the boundedness of the iterates from Theorem 5.1 and a recent idea in [lu2023unified, Theorem 3]. We provide its proof details in Appendix D.
@@
Let $f:\mathbb{R}^{n}\!\to\!\overline{\mathbb{R}}$ be a proper closed convex function with $S=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x)$ and $S\neq\emptyset$. The iterates 23 under $\text{A}^{\prime}$ with a positive sequence $\{c_{k}\}_{k\geq 0}$ bounded away from zero converge to $x^{\star}\in S$ asymptotically, and the cost value gaps satisfy
@@
Note that Theorems 5.2 and 5.1 can be viewed as the convergence counterpart for iPPM of Theorem 4.1 with two major differences: 1) Theorem 5.2 deals with the best iterate, unlike the last iterate in Theorem 4.1 (the guarantee for the average $\bar{x}_{k}=\frac{1}{k}\sum_{j=1}^{k}x_{j}$ or weighted average $\tilde{x}_{k}=({\sum_{j=0}^{k-1}c_{j}x_{j+1}})/({\sum_{j=0}^{k-1}c_{j}})$ is also straightforward; see remark 3 in the appendix); 2) the convergence of cost values in Theorem 5.2 relies on the boundedness of iterates in Theorem 5.1 whose proof is technically involved [rockafellar1976monotone, Theorem 1], while the convergence of iterates of exact PPM is established from the sublinear convergence of cost values in Theorem 4.1.
Similar to Theorem 4.2, linear convergence of iPPM can also be deduced when popular regularity conditions are assumed. However, the definitions of SC, RSI, EB, PL, and QG in Section 3 are stated over a sublevel set $f\leq f^{\star}+\nu]$ with $\nu>0$, which is convenient to estimate the iteration number to enter a certain region, such as the constant $k_{0}$ in Theorem 4.2; but those definitions become too stringent in the analysis. Indeed, note that the proof in Theorem 4.2 requires the next point $x_{k+1}$ remains inside the QG or PL region. As the stopping criteria A and B allow infeasible points, it is possible that $x_{k+1}$ is infeasible (i.e., $f(x_{k+1})=+\infty$), making the same proof fail. However, the iterates are still approaching an optimal solution and the inequality $\mu_{\mathrm{q}}/2\cdot\mathrm{dist}(x_{k+1},S)\leq f(x_{k+1})-f^{\star}$ with any $\mu_{\mathrm{q}}>0$ still holds automatically.
@@
Notice that we can also redefine RSI, EB, and PL using a neighborhood (similar to 24) and show the equivalence between them. However, we note that the neighborhood $\mathcal{U}$ for different regularity conditions may be different. For clarity, we present our final technical result that shows the linear convergence of iPPM under 24, which is the counterpart of Theorem 4.2.
@@
Let $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be a proper closed convex function with $S=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x)$ and $S\neq\emptyset$. Suppose $f$ satisfies 24 at all $x^{\star}\in S$. Let $\{x_{k}\}_{k\geq 0}$ be any sequence generated by iPPM 23 under A and B with a positive sequence $\{c_{k}\}_{k\geq 0}$ bounded away from zero. Then, there exist a nonnegative $\theta_{k}<1$ and a large $\bar{k}>0$ such that